On average, 10 patients visit a doctor’s surgery every hour. What is the probability that in the next 15 minutes, more than 1 patient will visit the surgery? Assuming the number of patients visiting a doctor’s surgery follows a Poisson distribution. a. 0.713 b. 0.918 c. 0.082 d. 0.205
Answer #1

Average = 10 patients every hour = (10/4) i.e. 2.5 patients per 15 minutes

Hence,

P(More than 1 patient)

= 1 - P(Less than or equal to 1)

= 1 - [\frac{e^{-2.5}2.5^0}{0!}+\frac{e^{-2.5}2.5^1}{1!}]

= 0.713

Option A is correct.